# Ratnangshu Das Robert Bosch Centre for Cyber-Physical Systems IISc, Bengaluru, India

### v1

@@ Abstract
This paper presents a Spatiotemporal Tube (STT)-based control framework for differential-drive mobile robots with dynamic uncertainties and external disturbances, guaranteeing the satisfaction of Temporal Reach-Avoid-Stay (T-RAS) specifications. The approach employs circular STT, characterized by smoothly time-varying center and radius, to define dynamic safe corridors that guide the robot from the start region to the goal while avoiding obstacles. In particular, we first develop a sampling-based synthesis algorithm to construct a feasible STT that satisfies the prescribed timing and safety constraints with formal guarantees. To ensure that the robot remains confined within this tube, we then design analytically a closed-form, approximation-free control law. The resulting controller is computationally efficient, robust to disturbances and model uncertainties, and requires no model approximations or online optimization. The proposed framework is validated through simulation studies on a differential-drive robot and benchmarked against state-of-the-art methods, demonstrating superior robustness, accuracy, and computational efficiency.
@@ 1 Introduction
Differential-drive mobile robots are among the most widely used robotic platforms due to their mechanical simplicity, maneuverability, and effectiveness in real-world navigation and manipulation tasks. They have been extensively adopted across diverse domains, including industrial applications [fragapane2022increasing], the healthcare sector [zhou2018fuzzy], autonomous exploration [ferreira2024autonomous], and many other fields. A fundamental challenge in deploying mobile robots is ensuring that they can reach designated targets in a fixed time while avoiding time-varying obstacles and adhering to state constraints. Temporal-reach-avoid-stay (T-RAS) specifications are crucial for addressing such challenges [Meng1]. T-RAS formulations also serve as building blocks for more complex temporal-logic tasks [Kloetzer, das2025spatiotemporal]. Therefore, developing and implementing safe and reliable control strategies is essential to perform these tasks effectively.
@@
A variety of control approaches have been developed to enforce reach-avoid-type specifications, including symbolic control, Control Barrier Functions (CBFs), and Model Predictive Control (MPC). Symbolic control techniques [tabuada2009verification] rely on abstractions of the state and input spaces and have shown effectiveness in solving complex temporal-logic tasks. However, these methods often suffer from severe computational overhead as the dimensionality or nonlinearity of the system increases, making them difficult to scale.
@@
To overcome discretization-related limitations, Control Barrier Functions (CBFs) [CBF] provide a continuous-state, optimization-based framework for synthesizing safety controllers. They have been successfully employed for dynamic obstacle avoidance [C3BF, tayal2024collision] and for enforcing temporal-logic constraints [CBF_STL, Meng3]. MPC [sun2017disturbance, jian2022dynamic] has also been widely employed to enforce reach-avoid and safety-critical specifications by optimizing future trajectories subject to dynamic and constraint models [MPC]. Despite solving the scalability issue of symbolic methods, CBF- and MPC-based control still relies on an optimization step, which can be computationally demanding for high-dimensional or fast-evolving systems. Moreover, they struggle to guarantee satisfaction of time-critical objectives such as Temporal reachability, and their performance degrades under dynamic uncertainties and external disturbances, which affect substantially real robots.
@@
In contrast, funnel-based control [PPC1] provides a natural framework capable of handling both external disturbances and time-critical objectives. It has been successfully applied to tracking control of unknown nonlinear systems [mishra2023approximation] and multi-agent coordination tasks [Funnel_STL_MAS]. However, the traditional funnel technique is not inherently equipped to address constraints such as obstacle avoidance. Some recent studies [RB, vrohidis2018prescribed] integrate articifial potential-field with funnel-invariance methods to establish collision-free navigation, without, however, accounting for dynamic uncertainties. The recently introduced KDF method [verginis2022kdf] combines sampling-based planning with funnel control to guarantee collision-free navigation for systems with uncertain dynamics, limited, however to fully actuated systems. To tackle that, KDF was recently extended in [lapandic2024kinodynamic] to account for underactuated surface vessels with uncertain dynamics, still limited to static environments. Additionally, [lapandic2024kinodynamic] restricts the control design to output only positive velocities, which, although suitable for surface vessels, can be conservative for general differential-drive mobile robots.
@@
To overcome the static-obstacle assumption, the spatiotemporal tube (STT) framework [STT] was introduced. STT define smooth, time-varying regions in space that evolve dynamically to form safe corridors around obstacles, allowing system trajectories to satisfy T-RAS objectives [das2025spatiotemporal]. In this work, we extend the Spatiotemporal Tube (STT) framework to a broader class of underactuated systems, focusing on differential-drive mobile robots, to ensure satisfaction of T-RAS specification. To achieve this, we introduce circular STT, characterized by smooth, time-varying center and radius, which serve as safe, time-evolving corridors guiding the robot toward its goal while avoiding obstacles. The contributions of this paper are twofold. First, we present a sampling-based synthesis procedure to construct circular STT that originate within the start region, strictly avoid obstacles, and reach the target region within a prescribed time horizon. Second, we derive a closed-form, approximation-free control law that guarantees the robot remains confined within the designed STT, thereby ensuring satisfaction of the T-RAS task. The proposed method is inherently robust to model uncertainties and external disturbances. The effectiveness of the proposed framework is demonstrated through simulation studies on a differential-drive robot, and its performance is benchmarked against state-of-the-art approaches, highlighting its accuracy, robustness, and computational efficiency.
@@ 2.1 Notation
For $a,b\in{\mathbb{N}}$ with $a\leq b$, we denote the closed interval in $\mathbb{N}$} as $a;b]:=\{a,a+1,\ldots,b\}$. A ball centered at $\mathsf{c}\in\mathbb{R}^{n}$ with radius $\mathsf{r}\in\mathbb{R}^{+}$ is defined as $\mathcal{B}(\mathsf{c},\mathsf{r}):=\{x\in\mathbb{R}^{n}\mid\|x-\mathsf{c}\|\leq\mathsf{r}\}$. The distance of a point $x\in{\mathbb{R}}^{n}$ from a set $A\subseteq{\mathbb{R}}^{n}$ is defined as $\text{dist}(x,A):=\min_{y\in A}\|x-y\|.$ All other notation in this paper follows standard mathematical conventions.
@@ 2.2 System Definition
$\mathcal{S}:\dot{\xi}=f(\xi,u)+d(\xi,t)\implies[\dot{x}_{1},\ \dot{x}_{2},\ \dot{\theta}]^{\top}=[v\cos(\theta),\ v\sin(\theta),\ \omega]^{\top}+d(\xi,t),$ (1)
@@ 2.3 System Specification
Let ${\mathbf{X}}\subseteq{\mathbb{R}}^{2}$ denote the state space and ${\mathbf{U}}:{\mathbb{R}}_{0}^{+}\rightarrow{\mathbf{X}}$ be the time-varying unsafe set. The initial and target sets are denoted by ${\mathbf{S}}\subset{\mathbf{X}}\setminus{\mathbf{U}}(0)$ and ${\mathbf{T}}\subset{\mathbf{X}}\setminus{\mathbf{U}}(t_{c})$, respectively, both assumed to be compact and connected. $t_{c}\in{\mathbb{R}}^{+}$ is the prescribed time of task completion. The control objective is formalized as a Temporal Reach-Avoid-Stay (T-RAS) task, defined below.
@@
Given the differential-drive system $\mathcal{S}$ in (1) and a T-RAS task in Definition (2.1), the goal is to design an approximation-free, closed form control law $v,\omega]^{\top}$, that ensures the system satisfies the specified T-RAS task.
@@ 3 Spatiotemporal Tube (STT)
To satisfy the given T-RAS specification, we leverage Spatiotemporal Tube (STT). STT act as time-varying structures in the state space that can effectively capture the dynamic constraints of a T-RAS task..
@@
$\Gamma(0)\subseteq{\mathbf{S}},\ \Gamma(t_{c})\subseteq{\mathbf{T}},\ \Gamma(t)\subseteq{\mathbf{X}}\setminus{\mathbf{U}}(t),\forall t\in[0,t_{c}],$ (2b)
@@
$\mathsf{c}_{i}(q_{\mathsf{c},i},t)$ $=\sum_{k=1}^{z_{i}}q_{\mathsf{c},i}^{(k)}b_{\mathsf{c},i}^{(k)}(t),i\in[1;2],\quad\mathsf{r}(q_{\mathsf{r}},t)=\sum_{k=1}^{z_{\mathsf{r}}}q_{\mathsf{r}}^{(k)}b_{\mathsf{r}}^{(k)}(t),$ (4)
@@
$\quad\|\mathsf{c}(q_{\mathsf{c}},0)-\mathsf{c}_{{\mathbf{S}}}\|=0,\ \mathsf{r}(c_{\mathsf{r}},0)=\mathsf{r}_{{\mathbf{S}}},\quad\|\mathsf{c}(q_{\mathsf{c}},t_{c})-\mathsf{c}_{{\mathbf{T}}}\|=0,\ \mathsf{r}(c_{\mathsf{r}},t_{c})=\mathsf{r}_{{\mathbf{T}}},$ (5a)
@@
$\quad\|\mathsf{c}(q_{\mathsf{c}},t)-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(c_{\mathsf{r}},t)-\mathsf{r}_{{\mathbf{X}}}\leq\eta,$ (5b)
$\quad-\mathsf{r}(q_{\mathsf{r}},t)+\mathsf{r}_{d}\leq\eta,$ (5c)
$\quad-\text{dist}(\mathsf{c}(q_{\mathsf{c}},t),{\mathbf{U}}(t))+\mathsf{r}(q_{\mathsf{r}},t)\leq\eta,$ (5d)
where ${\mathbb{B}}(\mathsf{c}_{\mathbf{S}},\mathsf{r}_{\mathbf{S}})\subseteq{\mathbf{S}}$, ${\mathbb{B}}(\mathsf{c}_{{\mathbf{T}}},\mathsf{r}_{\mathbf{T}})\subseteq{\mathbf{T}}$, and ${\mathbb{B}}(\mathsf{c}_{{\mathbf{X}}},\mathsf{r}_{\mathbf{X}})\subseteq{\mathbf{X}}$. $\mathsf{r}_{d}\in{\mathbb{R}}^{+}$ is a user-defined lower bound on the tube radius.
One can readily observe that a solution to the ROP with $\eta^{*}\leq 0$ ensures that all conditions in Definition 3.1 are satisfied.
@@
$\quad\|\mathsf{c}(q_{\mathsf{c}},0)-\mathsf{c}_{{\mathbf{S}}}\|=0,\ \mathsf{r}(c_{\mathsf{r}},0)=\mathsf{r}_{{\mathbf{S}}},\quad\|\mathsf{c}(q_{\mathsf{c}},t_{c})-\mathsf{c}_{{\mathbf{T}}}\|=0,\ \mathsf{r}(c_{\mathsf{r}},t_{c})=\mathsf{r}_{{\mathbf{T}}},$ (7a)
@@
$\quad\|\mathsf{c}(q_{\mathsf{c}},t_{s})-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(c_{\mathsf{r}},t_{s})-\mathsf{r}_{{\mathbf{X}}}\leq\eta,$ (7b)
$\quad-\mathsf{r}(q_{\mathsf{r}},t_{s})+\mathsf{r}_{d}\leq\eta,$ (7c)
$\quad-\text{dist}(\mathsf{c}(q_{\mathsf{c}},t_{s}),{\mathbf{U}}(t))+\mathsf{r}(q_{\mathsf{r}},t)\leq\eta,$ (7d)
@@
holds, where $\mathcal{L}=\mathcal{L}_{\mathsf{c}}+\mathcal{L}_{\mathsf{r}}$, then the resulting STT
@@
(i) $-\mathsf{r}(q_{\mathsf{r}}^{*},t)+\mathsf{r}_{d}=-\mathsf{r}(q_{\mathsf{r}}^{*},t_{s})+\mathsf{r}_{d}+\mathsf{r}(q_{\mathsf{r}}^{*},t_{s})-\mathsf{r}(q_{\mathsf{r}}^{*},t)\leq\ \eta^{*}+\mathcal{L}_{\mathsf{r}}|t-t_{s}|\leq\eta^{*}+\mathcal{L}_{\mathsf{r}}\varepsilon\leq\eta^{*}+\mathcal{L}\epsilon\leq 0.$
@@
(ii) $\|\mathsf{c}(q_{\mathsf{c}},t)-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(c_{\mathsf{r}},t)-\mathsf{r}_{{\mathbf{X}}}=\|\mathsf{c}(q_{\mathsf{c}},t)-\mathsf{c}_{{\mathbf{X}}}\|-\|\mathsf{c}(q_{\mathsf{c}},t_{s})-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(c_{\mathsf{r}},t)-\mathsf{r}(c_{\mathsf{r}},t_{s})$
$\hskip 85.35826pt+\|\mathsf{c}(q_{\mathsf{c}},t_{s})-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(c_{\mathsf{r}},t_{s})-\mathsf{r}_{{\mathbf{X}}}\leq\mathcal{L}_{\mathsf{c}}|t-t_{s}|+\mathcal{L}_{\mathsf{r}}|t-t_{s}|+\eta^{*}\leq\eta^{*}+\mathcal{L}\epsilon\leq 0.$
@@
$=$ $-\text{dist}(\mathsf{c}(q_{\mathsf{c}},t),{\mathbf{U}}(t))+\text{dist}(\mathsf{c}(q_{\mathsf{c}},t_{s}),{\mathbf{U}}(t_{s}))+\mathsf{r}(q_{\mathsf{r}},t)-\mathsf{r}(q_{\mathsf{r}},t_{s})-\text{dist}(\mathsf{c}(q_{\mathsf{c}},t_{s}),{\mathbf{U}}(t_{s}))+\mathsf{r}(q_{\mathsf{r}},t_{s})$
$\leq$ $\|\mathsf{c}(q_{\mathsf{c}},t)-\mathsf{c}(q_{\mathsf{c}},t_{s})\|+\|\mathsf{r}(q_{\mathsf{r}},t)-\mathsf{r}(q_{\mathsf{r}},t_{s})\|+\eta_{S}^{*}\leq(\mathcal{L}_{\mathsf{c}}+\mathcal{L}_{\mathsf{r}})\epsilon\leq\mathcal{L}\epsilon+\eta_{S}^{*}\leq 0$
@@
The Lipschitz constants $\mathcal{L}_{\mathsf{r}}$ and $\mathcal{L}_{\mathsf{c}}$ are needed to verify condition (8). An estimation procedure for these constants is provided in [das2025spatiotemporal].

### v2

@@ Abstract
This paper presents a computationally lightweight and robust control framework for differential-drive mobile robots with dynamic uncertainties and external disturbances, guaranteeing the satisfaction of Temporal Reach-Avoid-Stay (T-RAS) specifications. The approach employs circular spatiotemporal tubes (STTs), characterized by smoothly time-varying center and radius, to define dynamic safe corridors that guide the robot from the start region to the goal while avoiding obstacles. In particular, we first develop a sampling-based synthesis algorithm to construct a feasible STT that satisfies the prescribed timing and safety constraints with formal guarantees. To ensure that the robot remains confined within this tube, we then analytically design a closed-form control that is computationally efficient and robust to disturbances. The proposed framework is validated through simulation studies on a differential-drive robot and benchmarked against state-of-the-art methods, demonstrating superior robustness, accuracy, and computational efficiency.
@@ 1 Introduction
Differential-drive mobile robots are among the most widely used robotic platforms due to their mechanical simplicity, maneuverability, and effectiveness in real-world navigation and manipulation tasks. They have been extensively adopted across diverse domains, including industrial applications [10], the healthcare sector [28], autonomous exploration [9], and many other fields. A fundamental challenge in deploying mobile robots is ensuring that they can reach designated targets in a fixed time while avoiding time-varying obstacles and adhering to state constraints. Temporal-reach-avoid-stay (T-RAS) specifications are crucial for addressing such challenges [18]. T-RAS formulations also serve as building blocks for more complex temporal-logic tasks [14, 6]. Therefore, developing and implementing safe and reliable control strategies is essential to perform these tasks effectively.
@@
A variety of control approaches have been developed to enforce reach-avoid-type specifications, including symbolic control, Control Barrier Functions (CBFs), and Model Predictive Control (MPC). Symbolic control techniques [23] rely on abstractions of the state and input spaces and have shown effectiveness in solving complex temporal-logic tasks. However, these methods often suffer from severe computational overhead as the dimensionality or nonlinearity of the system increases, making them difficult to scale.
@@
To overcome discretization-related limitations, Control Barrier Functions (CBFs) [2] provide a continuous-state, optimization-based framework for synthesizing safety controllers. They have been successfully employed for dynamic obstacle avoidance [11, 24] and for enforcing temporal-logic constraints [16, 17]. MPC [22, 13] has also been widely employed to enforce reach-avoid and safety-critical specifications by optimizing future trajectories subject to dynamic and constraint models [12]. Despite solving the scalability issue of symbolic methods, CBF- and MPC-based control still relies on an optimization step, which can be computationally demanding for high-dimensional or fast-evolving systems. Moreover, they struggle to guarantee satisfaction of time-critical objectives such as Temporal reachability, and their performance degrades under dynamic uncertainties and external disturbances, which affect substantially real robots.
@@
In contrast, funnel-based control [3] provides a natural framework capable of handling both external disturbances and time-critical objectives. It has been successfully applied to tracking control of unknown nonlinear systems [19] and multi-agent coordination tasks [4]. However, the traditional funnel technique is not inherently equipped to address constraints such as obstacle avoidance. Some recent studies [20, 27] integrate articifial potential-field with funnel-invariance methods to establish collision-free navigation, without, however, accounting for dynamic uncertainties. The recently introduced KDF method [26] combines sampling-based planning with funnel control to guarantee collision-free navigation for systems with uncertain dynamics, limited, however to fully actuated systems. To tackle that, KDF was recently extended in [15] to account for underactuated surface vessels with uncertain dynamics, still limited to static environments. Additionally, [15] restricts the control design to output only positive velocities, which, although suitable for surface vessels, can be conservative for general differential-drive mobile robots.
@@
To overcome the static-obstacle assumption, the spatiotemporal tube (STT) framework [7] was introduced. STT define smooth, time-varying regions in space that evolve dynamically to form safe corridors around obstacles, allowing system trajectories to satisfy T-RAS objectives [6]. In this work, we extend the Spatiotemporal Tube (STT) framework to a broader class of systems, specifically the underactuated systems, with the main focus lying on differential-drive mobile robots, to ensure satisfaction of T-RAS specification. To achieve this, we introduce circular STT, characterized by smooth, time-varying center and radius, which serve as safe, time-evolving corridors guiding the robot toward its goal while avoiding obstacles. The contributions of this paper are twofold. First, we present a sampling-based synthesis procedure to construct circular STT that originate within the start region, strictly avoid obstacles, and reach the target region within a prescribed time horizon. Second, we derive a closed-form control law that is inherently robust to external disturbances and guarantees the robot remains confined within the designed STT, thereby ensuring satisfaction of the T-RAS task. The effectiveness of the proposed framework is demonstrated through simulation studies on a differential-drive robot, and its performance is benchmarked against state-of-the-art approaches, highlighting its accuracy, robustness, and computational efficiency.
@@ 2.1 Notation
For $a,b\in{\mathbb{N}}$ with $a\leq b$, we denote the closed interval in $\mathbb{N}$} as $a;b]:=\{a,a+1,\ldots,b\}$. A ball centered at $\mathsf{c}\in\mathbb{R}^{n}$ with radius $\mathsf{r}\in\mathbb{R}^{+}$ is defined as $\mathcal{B}(\mathsf{c},\mathsf{r}):=\{x\in\mathbb{R}^{n}\mid\|x-\mathsf{c}\|\leq\mathsf{r}\}$. The distance of a point $x\in{\mathbb{R}}^{n}$ from a set $A\subseteq{\mathbb{R}}^{n}$ is defined as $\text{dist}(x,A):=\inf_{y\in A}\|x-y\|.$ All other notation in this paper follows standard mathematical conventions.
@@ 2.2 System Definition
$\mathcal{S}:\dot{\xi}=f(\xi,u)+d(\xi,t)\implies\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\\ \dot{\theta}\end{bmatrix}=\begin{bmatrix}v\cos(\theta)\\ v\sin(\theta)\\ \omega\end{bmatrix}+d(\xi,t),$ (1)
@@ 2.3 System Specification
Let ${\mathbf{X}}\subseteq{\mathbb{R}}^{2}$ denote the state space and ${\mathbf{U}}:{\mathbb{R}}_{0}^{+}\rightarrow{\mathbf{X}}$ be the time-varying unsafe set, whose maximum permissible velocity is given by $\mathcal{L}_{U}$. In this work, the time evolution of the unsafe set is considered to be known a priori, but that does not limit the scope of the work. Such an assumption is standard in<§ the literature of reach-avoid problems [1, 5]. The initial and target sets are denoted by ${\mathbf{S}}\subset{\mathbf{X}}\setminus{\mathbf{U}}(0)$ and ${\mathbf{T}}\subset{\mathbf{X}}\setminus{\mathbf{U}}(t_{c})$, respectively, both assumed to be compact and connected. $t_{c}\in{\mathbb{R}}^{+}$ is the prescribed time of task completion. The control objective is formalized as a Temporal Reach-Avoid-Stay (T-RAS) task, defined below.
@@
Given the differential-drive system $\mathcal{S}$ in (1) and a T-RAS task in Definition (2.1), the goal is to design a closed form control law $v,\omega]^{\top}$, that is robust to external disturbances and ensures the system satisfies the specified T-RAS task.
@@ 3 Spatiotemporal Tube (STT)
To satisfy the given T-RAS specification, we leverage Spatiotemporal Tube (STT). STT act as time-varying structures in the state space that can effectively capture the dynamic constraints of a T-RAS task.
@@
$\Gamma(0)\subseteq{\mathbf{S}},\ \Gamma(t_{c})\subseteq{\mathbf{T}},\ \Gamma(t)\subseteq{\mathbf{X}}\setminus{\mathbf{U}}(t),\forall t\in[0,t_{c}].$ (2b)
@@
$\mathsf{c}_{i}(q_{\mathsf{c},i},t)$ $=\sum_{k=1}^{z_{i}}q_{\mathsf{c},i}^{(k)}b_{\mathsf{c},i}^{(k)}(t),i\in[1;2],$ (4)
$\mathsf{r}(q_{\mathsf{r}},t)$ $=\sum_{k=1}^{z_{\mathsf{r}}}q_{\mathsf{r}}^{(k)}b_{\mathsf{r}}^{(k)}(t),$ (5)
@@
$\quad\|\mathsf{c}(q_{\mathsf{c}},0)-\mathsf{c}_{{\mathbf{S}}}\|=0,\ \mathsf{r}(c_{\mathsf{r}},0)=\mathsf{r}_{{\mathbf{S}}},$ (6a)
$\quad\|\mathsf{c}(q_{\mathsf{c}},t_{c})-\mathsf{c}_{{\mathbf{T}}}\|=0,\ \mathsf{r}(c_{\mathsf{r}},t_{c})=\mathsf{r}_{{\mathbf{T}}},$ (6b)
@@
$\quad\|\mathsf{c}(q_{\mathsf{c}},t)-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(c_{\mathsf{r}},t)-\mathsf{r}_{{\mathbf{X}}}\leq\eta,$ (6c)
$\quad-\mathsf{r}(q_{\mathsf{r}},t)+\mathsf{r}_{d}\leq\eta,$ (6d)
$\quad-\text{dist}(\mathsf{c}(q_{\mathsf{c}},t),{\mathbf{U}}(t))+\mathsf{r}(q_{\mathsf{r}},t)\leq\eta,$ (6e)
where ${\mathbb{B}}(\mathsf{c}_{\mathbf{S}},\mathsf{r}_{\mathbf{S}})\subseteq{\mathbf{S}}$, ${\mathbb{B}}(\mathsf{c}_{{\mathbf{T}}},\mathsf{r}_{\mathbf{T}})\subseteq{\mathbf{T}}$, and ${\mathbb{B}}(\mathsf{c}_{{\mathbf{X}}},\mathsf{r}_{\mathbf{X}})\subseteq{\mathbf{X}}$, with $\mathsf{c}_{\mathbf{S}}\in{\mathbf{S}},\mathsf{c}_{\mathbf{T}}\in{\mathbf{T}},\mathsf{c}_{\mathbf{X}}\in{\mathbf{X}}$ and $\mathsf{r}_{\mathbf{S}},\mathsf{r}_{\mathbf{T}},\mathsf{r}_{\mathbf{X}}\in{\mathbb{R}}^{+}.$ $\mathsf{r}_{d}\in{\mathbb{R}}^{+}$ is a user-defined lower bound on the tube radius.
Note that $\eta$ is introduced as a slack variable, which has to be minimized in order to reduce the worst-case violations of the constraints across the dimensions. One can readily observe that a solution to the ROP with $\eta^{*}\leq 0$ ensures that all conditions in Definition 3.1 are satisfied. The proposed solution is sound: whenever a feasible solution exists, the resulting tubes guarantee that the robot will reach the goal, avoiding the obstacles. The feasibility of the optimization problem (6), however, depends on the degree of the polynomial basis functions used to parameterize the STTs. If no feasible solution is obtained, then a higher degree polynomial may be required to represent the tubes.
@@
$\quad\|\mathsf{c}(q_{\mathsf{c}},0)-\mathsf{c}_{{\mathbf{S}}}\|=0,\ \mathsf{r}(q_{\mathsf{r}},0)=\mathsf{r}_{{\mathbf{S}}},$ (8a)
$\quad\|\mathsf{c}(q_{\mathsf{c}},t_{c})-\mathsf{c}_{{\mathbf{T}}}\|=0,\ \mathsf{r}(q_{\mathsf{r}},t_{c})=\mathsf{r}_{{\mathbf{T}}},$ (8b)
@@
$\quad\|\mathsf{c}(q_{\mathsf{c}},t_{s})-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(q_{\mathsf{r}},t_{s})-\mathsf{r}_{{\mathbf{X}}}\leq\eta,$ (8c)
$\quad-\mathsf{r}(q_{\mathsf{r}},t_{s})+\mathsf{r}_{d}\leq\eta,$ (8d)
$\quad-\text{dist}(\mathsf{c}(q_{\mathsf{c}},t_{s}),{\mathbf{U}}(t))+\mathsf{r}(q_{\mathsf{r}},t_{s})\leq\eta.$ (8e)
@@
holds, where $\mathcal{L}=\mathcal{L}_{\mathsf{c}}+\mathcal{L}_{U}+\mathcal{L}_{\mathsf{r}}$, then the resulting STT
@@
(i) $-\mathsf{r}(q_{\mathsf{r}}^{*},t)+\mathsf{r}_{d}=-\mathsf{r}(q_{\mathsf{r}}^{*},t_{s})+\mathsf{r}_{d}+\mathsf{r}(q_{\mathsf{r}}^{*},t_{s})-\mathsf{r}(q_{\mathsf{r}}^{*},t)$
$\leq\ \eta^{*}+\mathcal{L}_{\mathsf{r}}|t-t_{s}|\leq\eta^{*}+\mathcal{L}_{\mathsf{r}}\varepsilon\leq\eta^{*}+\mathcal{L}\epsilon\leq 0.$
@@
(ii) $\|\mathsf{c}(q_{\mathsf{c}},t)-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(c_{\mathsf{r}},t)-\mathsf{r}_{{\mathbf{X}}}$
$=$ $\|\mathsf{c}(q_{\mathsf{c}},t)-\mathsf{c}_{{\mathbf{X}}}\|-\|\mathsf{c}(q_{\mathsf{c}},t_{s})-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(c_{\mathsf{r}},t)-\mathsf{r}(c_{\mathsf{r}},t_{s})+\|\mathsf{c}(q_{\mathsf{c}},t_{s})-\mathsf{c}_{{\mathbf{X}}}\|+\mathsf{r}(c_{\mathsf{r}},t_{s})-\mathsf{r}_{{\mathbf{X}}}$
$\leq$ $\mathcal{L}_{\mathsf{c}}|t-t_{s}|+\mathcal{L}_{\mathsf{r}}|t-t_{s}|+\eta^{*}\leq\eta^{*}+\mathcal{L}\epsilon\leq 0.$
@@
$=$ $-\text{dist}(\mathsf{c}(q_{\mathsf{c}},t),{\mathbf{U}}(t))+\text{dist}(\mathsf{c}(q_{\mathsf{c}},t_{s}),{\mathbf{U}}(t))$
$-\text{dist}(\mathsf{c}(q_{\mathsf{c}},t_{s}),{\mathbf{U}}(t))+\text{dist}(\mathsf{c}(q_{\mathsf{c}},t_{s}),{\mathbf{U}}(t_{s}))$
$+\mathsf{r}(q_{\mathsf{r}},t)-\mathsf{r}(q_{\mathsf{r}},t_{s})-\text{dist}(\mathsf{c}(q_{\mathsf{c}},t_{s}),{\mathbf{U}}(t_{s}))+\mathsf{r}(q_{\mathsf{r}},t_{s})$
$\leq$ $\|\mathsf{c}(q_{\mathsf{c}},t)-\mathsf{c}(q_{\mathsf{c}},t_{s})\|+\mathcal{L}_{U}|t-t_{s}|+\|\mathsf{r}(q_{\mathsf{r}},t)-\mathsf{r}(q_{\mathsf{r}},t_{s})\|+\eta_{S}^{*}\leq(\mathcal{L}_{\mathsf{c}}+\mathcal{L}_{U}+\mathcal{L}_{\mathsf{r}})\epsilon\leq\mathcal{L}\epsilon+\eta_{S}^{*}\leq 0.$
@@
The Lipschitz constants $\mathcal{L}_{\mathsf{r}}$ and $\mathcal{L}_{\mathsf{c}}$ are needed to verify condition (9). An estimation procedure for these constants is provided in [6].